OpenStudy (anonymous):

What should be added to both sides of this equation to solve for the variable? p - 33 = 17

OpenStudy (anonymous):

33

OpenStudy (anonymous):

add 33 to both sides.

OpenStudy (anonymous):

p -33 +33 = 17 + 33 p = 50

OpenStudy (anonymous):

p -33 + 33 = 17 + 33 p = 17 + 33 p = 50